Items Values

Indirect Business taxes and subsidies 416.2

Corporate Profit    1,239.5

Personal tax Rate    14.6461%

Personal Consumption Expenditure 3,470.8

Capital Consumption Allowance 10.55357%

Investing spending    774.3

Export 335.12

Transfer Payment    1,402.4

Gov. purchase of goods    1,036.0

Imports    383.82

"What is the" Net national product ?

" " Personal Income ?

" "National income?

" "Disposable personal income ?

I'm not sure why my teacher has that long percentage for Capital Consumption Allowance. Also, I read where it can be used for Depreciation too. Maybe he's trying to trick us.

I did the GDP C+I+G+(Xn)

3,470.8+774.3+1036+335.12-383.82=5232.4
Answer #1

(a) Net national product is = Gross national product - Depreciation,

Depreciation is not given here, so here NNP = GNP

GNP = Y = C + I + G + X + Z

where C = Consumption

I = Investment

G =  Government

X = net exports

Z = net income earned by domestic residents from overseas investments - net income earned by foreign residents from domestic investments

GNP = 3,470.8 + 774.3 + 1036 + 335.12 - 383.82

GNP = 5232.4

(b) Personal Income = National income - social security contribution tax - corporate income tax - retained earnings + transfer payments

here corporate income tax would be 14.6461% of 1239.5 = $181.53

Personal income = 5232.4 - $181.53 + 1402.4 = $6453.27

(c) National income = National domestic product - Indirect tax  

= 5232.4 - 416.2

= $4816.2

(d) Disposable income = Personal taxes - indirect taxes

= 6453.27 - 416.2

= 5992.07